The measure of angle A is 30 degrees less than twice the measures of angle B. Find the measures of each angle of the parallelogram
Answer by stanbon(75887)

The measure of angle A is 30 degrees less than twice the measures of angle B. Find the measures of each angle of the parallelogram
The measure of angle A is 30 degrees less than twice the measures of angle B. Find the measures of each angle of the parallelogram
---
I'll assume that A and B are consecutive angles of the parallelogram.
A and B are suppliemntary.
A + B = 180
A = 2B-30
--------------
2B-30+B = 180
3B = 210
B = 70 degrees
----
A = 180-70 = 110 degrees
------------------------------
Cheers,
Stan H.
-------------